# StyleBooth: Image Style Editing with Multimodal Instruction

### v1

@@ Abstract
Keywords:
Text-based style editing Exemplar-based style editing Multimodal instruction-tuning
@@ 1 Introduction
To tackle these challenges, we introduce StyleBooth, a unified method for text and exemplar-based style editing that accommodates hybrid editing instructions combining written text and exemplar images. Concretely, we devise a unified conditioning schema for the diffusion model, using the tokens "<style>" and "<image>" as identifiers to indicate arbitrary style and exemplar style images. The text instruction, which substitutes the "<style>" token with specified style words, is processed through the text encoder. Additionally, the exemplar image is encoded by the image encoder, and the image feature is aligned with the encoded text feature in a unified hidden space, forming the multimodal input for the diffusion model. Merging the flexibility of natural language with the explicitness of exemplar images enables the exploration of more creative and personalized styles, such as style interpolation and style composition. We present StyleBooth models learning from content-strong-aligned and high-resolution image pairs, which is capable of generating high content fidelity and quality edited images. Samples of edited results are shown in Fig. 1. To support StyleBooth, we also present a high-quality style editing dataset, constructed through an effective pipeline combined iterative edit tuning and usability filtering. This strategy is easy to generalize to any other image editing tasks.
@@ 2.1 Image Style Transfer
Image style transfer has been widely studied by researchers. It involves the process of applying stylistic representations to content images in order to generate images with specific styles. Style transfer based on deep neural networks  [13, 14] typically employs separating and recombining the content and style of an image to achieve a simple transfer of style, but there is a problem with the precision of the style representation. There are also works based on GANs  [22, 53, 9, 26] that, through training generators and discriminators, apply the representations of style images while preserving the features of content images as much as possible. However, these methods lack sufficient adaptability and the trained models need to correspond to the style, which limits the possibility of more extensive transformation. Recently, diffusion-based image generation has developed rapidly and has also been applied to style transfer tasks. These methods [52, 8, 43, 47] often require users to provide only a few or even just one style image to achieve a relatively universal style transfer. In this process, generating high-quality transferred images while maintaining the semantic information of the content image remains a challenge. Our StyleBooth builds upon this foundation and can continuously generate high-quality transferred images while preserving the semantic information of the content image.
2.2 Text-Based Image Editing
@@
Image editing is a common application in practical scenarios, where it can achieve customized image generation through external guidance. Many pre-trained generative models based on diffusion models [18, 45, 46, 34, 38] can generate corresponding images according to text, but it is difficult to finely edit images according to specific needs while keeping the main content unchanged. Recent works edit global semantic information of images based on text guidance [3, 16, 32]. However, they still require detailed descriptions or corrections of the target images. Text-based image editing offers great convenience, as it only requires describing the specific changes to be made to the image, such as "Let it be in the style of Picasso". InstructPix2Pix [4] is trained with instructional data generated by GPT-3 [5] and Prompt-to-Prompt [16]. MagicBrush [48] builds an image editing dataset with manually annotated instructions using online image editing tools. Emu Edit [42] proposes a multi-task image editing model and trains it on image editing and recognition data. Although these works have achieved improvements, high-quality data and precise execution of editing instructions remain significant challenges.
2.3 Exemplar-Based Style Editing
@@
Exemplar-based image editing typically uses one or several image exemplars as references for generations. InST [52] introduces an attention-based textual inversion method that inverts a painting into corresponding text embeddings to guide the model in creating images with a specific artistic appearance. StyleDrop [43] employs an adapter fine-tuning approach to efficiently train models corresponding to the reference images. RAVIL [51] modifies the denoising inference chain of diffusion models to align with the real image inversion chain, generating diversified and high-quality variants of real-world images. The aforementioned methods only support editing with exemplar images, but for more detailed guidance, our StyleBooth can generate highly creative images based on any image exemplar and text instruction while maintaining the consistency of the identity of the images to be edited.
@@
3 Method
In this section, we present a unified text and exemplar-based style editing method namely StyleBooth. For styles easy to be explained in text, the text-based method is capable of following text instructions. As for those hard to write accurately, one may use an exemplar-based method providing the target style by uploading a reference image. StyleBooth has both advantages.
High-quality data is also essential to style editing tasks. We introduce a paired image constructing strategy based on iterative style-destyle editing and usability control. Following this methodology, we present a high-quality dataset supporting text and exemplar-based style editing.
3.1 StyleBooth
StyleBooth unifies text instruction and exemplar image into a multimodal instruction. Our model is a latent diffusion model conditioned by this multimodal instruction together with an image to be edited as illustrated in Fig. 2.
3.1.1 Preliminaries.
Diffusion models[18] are trained to sample image $x$ from Gaussian noise through a reversed diffusion process, which is a sequence of denoising operations of predicting the added noise $\epsilon\sim\mathcal{N}(0,1)$ at a specific timestamp $t$. To improve the efficiency, latent diffusion[38] utilizes a pre-trained variational autoencoder[27] $\mathcal{E}$ to encode $x$ into latent space $z=\mathcal{E}(x)$ and operates denoising in the compressed latent space.
InstructPix2Pix[4], an image editing model, tuned from a pre-trained T2I latent diffusion model $\epsilon_{\theta}$ conditioned by instruction text $c_{T}$ and original image $c_{I}$. $c_{T}$ is encoded by pre-trained CLIP[36] text encoder into text features $h_{T}=\mathcal{C}_{T}(c_{T})$, while $c_{I}$ is encoded by a pre-trained variational autoencoder into latent. The training objective is formulated as follows:
@@ 3.1 StyleBooth
3.1.2 Multimodal Instruction.
To support text-based and exemplar-based style editing at the same time, our model supports text-only, image-only, and text-image instructions, as illustrated in Fig. 2. Input images are treated as an undetermined identifier "<image>" in the text encoder.
We encode the style exemplar image $c_{E}$ using CLIP image encoder $\mathcal{C}_{I}$. The patch features after the last Transformer layer are considered as exemplar image features. We introduce a trainable convolutional alignment layer $W$ to map image features into text feature space:
@@
3.1.3 Textual Instruction Templates for Multimodal Inputs.
For diverse, comprehensible, and grammar-correct style editing instructions, we leverage the Turbo 128k version of GPT4[35] to generate editing instructions. Instead of complete sentences, we ask LLM for literal templates composed of normal text and identifier tokens. These identifiers represent the undetermined target styles. We write examples to make sure LLM to precisely interpret our intent. For text-based style editing, we give an instruction example "Let this image be in the style of <style>". For exemplar-based style editing, we use "Let this image be in the style of <image>" as an instruction example. Notice that we use "<style>" and "<image>" as the identifiers of arbitrary styles in types of text and exemplar. Adequate templates are generated for each task for training. During inference, input instructions can be varied and unnecessarily limited to the training instructions.
3.1.4 Text-based and Exemplar-Based Style Editing.
Text-based methods perform editing following text-only instructions. During training, instruction templates are randomly selected and the identifier "<style>" is replaced by the actual style name. We follow the methodology of InstructPix2Pix and fine-tune from their pre-trained model. Supervised by the paired images, the model learns to add the specific style following text-only instructions.
Style of exemplar image and that of target image must be well-aligned during training phase. We search for the most appropriate exemplars from other target images in the same style subset. Within a style subset, we first calculate CLIP score for every target image with other in-class stylized images. Images with the highest score are the most similar ones and they are the most appropriate candidates as exemplar-style images. During training, the final exemplar will be randomly selected from the top 10 most similar images.
An instruction template is randomly selected and encoded into text features by $\mathcal{C_{T}}$, where the identifier "<image>" is treated as an unknown identifier at the beginning. Composed of exemplar image features, the original textual instruction is transformed into a multimodal instruction $h$. The training objective is changed into:
@@
3.1.5 Scale Weighting Mechanism for Compositional Style Editing.
As shown in Fig. 2, in order to balance the presence of visual elements, we introduce a Scale Weighting Mechanism. For $n$ textual or visual style representations, we assign a series of multiplicative scale factors $\alpha_{1},\alpha_{2},\ldots,\alpha_{n}$. We find the Scale Weighting Mechanism effective in controlling the style presence ratio in the final results.
@@ 3.2 High-Quality Style Editing Dataset
Intuitively, images in specific styles can be generated by any T2I model as long as the style-related keywords to the text prompt. We follow the same route but design text prompts carefully through prompt expansion provided by Fooocus[11] to avoid the style of generated image misaligned with our prompts. For each of the styles, prompts are modified by the style format before being sent to T2I model. We accept 67 prompt formats and 217 diverse content prompts, ending up with 67 different styles and 217 images per style. This batch of stylized images is marked as $BatchA$. Next we generate $BatchB$ consisting of plain images. For diversity and quality of generated images, we randomly select image captions from the LAION Art[41] dataset as prompts. Unlike style image generation, no specific style prompt expansion is used.
3.2.2 Vanilla De-Style and Usability Filtering.
@@
We train a vanilla version of image editing model following the methodology of Instruct-Pix2Pix[4]. This model has a basic I2I editing ability but the generated data quality falls far short of our needs. Starting from $BatchA$, we use the image editing model to de-style them into plain photographic images generating the corresponding $BatchA^{1}$. As shown in the second column of Fig. 4, after the first de-style operation, the content structure is preserved yet most of these de-styled images contain features of the original style. We then employ a CLIP-based[36] metric to filter out image pairs that are too similar, indicating no significant style change, and that are overly different, which could imply potential content distortion. By establishing upper and lower thresholds for CLIP scores, this filtering process effectively eliminates the majority of evident failures produced by the image editing model.
Figure 4: Generated samples of the intermediate and final image pairs during Iterative Style-Destyle Tuning and Editing. As the iteration process, image quality gets higher while key style features are gradually wiped off in the de-styled images. We show the style images and de-style results generated in 1st and 2nd de-styled phase and a plain image and results generated in 1st style phase.
3.2.3 Iterative Style-Destyle Tuning and Editing.
@@
To further strengthen the pre-trained I2I model, we take filtered image pairs as training data and fine-tune it using Style Tuner to learn style-adding ability. Now, the $BatchA^{1}$ are used as inputs while the style images in $BatchA$ as targets. Efficient tuning methods[20, 24, 15, 25] show strong few-shot learning capability. Here, we add efficient tuners called Style Tuners to the basic I2I model and tune different styles separately which means each tuner is only responsible for one specific style. After tuning, the base model with tuners will edit $BatchB$ images to add the styles they have learned. Then we get $BatchB^{1}$. Some of the results are shown in the last column of Fig. 4. The style-adding results is frustrating and styles in $BatchB^{1}$ is still not that accurate. Again, $BatchB$ and $BatchB^{1}$ images are measured and filtered by CLIP-based metric to be the next round training data. Note that we always use the high-quality images in $BatchA$ or $BatchB$ as the training targets. In this round, we tune I2I model to be a style removal model using De-style Tuner, which means images in $BatchB^{1}$ are used as inputs and image in $BatchB$ as targets. Still, one tuner is trained to remove one type of style. The tuned I2I model edits the stylized image in $BatchA$ and geneates $BatchA^{2}$ which are shown in the third row of Fig. 4. The image quality in $BatchA^{2}$ is improved dramatically. Then CLIP scores are calculated to filter out the unexpected image pairs.
After several iterations, the $BatchA^{n}$ are generated and image pairs of $BatchA$ and $BatchA^{n}$ are filtered by CLIP. The style-diverse, high-quality and content-preserved stylized/plain image pairs are finally observed. We compare the number of images those passing the CLIP filter. Comparing to the results of vanilla de-style, the average usability rate is increased dramatically from 38.11% to 79.91% by 41.80%, which shows that our Iterative Style-Destyle Editing is effective for quality improvement of paired images.
@@ 4.1 Experimental setup
We conduct our data generation and experiments on SCEPTER[40], an open-
source project for image generation, which not only integrates plenty of styles from Fooocus[11] but also provides an easy implementation of efficient tuning.
For generating BatchA and BatchB images, we use an advanced T2I diffusion model SDXL[1]. We accept LoRAs[20] with rank 256 as the implementation of Style Tuners and De-style Tuners. The editing-tuning iterations stop at the end of the second de-style process. The pairs of style and de-styled images are the final training pairs.
The trainable matrix $W$ is configured as a $6\times 6$ convolutional layer with a stride of 4, which maps a $14\times 14$ grid of visual features into a space of 9 textual features. We utilize 8 NVIDIA A800-SXM4-80GB GPUs for the training and validation processes. During inference, we implement classifier-free guidance for both image and text conditions, akin to the approach described in [4]. Following the recommendation in [28], we also apply a re-scaling factor of 0.5 to the outcomes. For instruction-based style editing, we fine-tune the pre-trained model of [4] for 5000 steps under the learning rate of 0.0001. For exemplar-based style editing, we only tune the convolutional alignment layers and the decoder of the pre-trained U-Net[39] under the learning rate of 0.00001 for 35000 steps. Both are trained using Adam[30] optimizer. The scale weighting is only applied during inference in compositional style editing tasks. We adjust scale factors in the range of $[0.5,1.5]$.
4.2 Text-Based Style Editing
4.2.1 Baselines and Benchmarks.
We compare our model with 3 state-of-the-art text-based image editing methods. InstructPix2Pix[4] presents a machine-generated image editing instruction-tuning dataset involving triplets of the original image, edited image, and text instruction. Finetuned using this dataset, the T2I (Text-to-Image) model demonstrates proficiency in executing image editing tasks, encompassing style alterations. Magic Brush[48] introduces a multi-turn editing task and tunes the model in human-annotated local inpainting image pairs, mainly focusing on the local editing task. The Emu Edit[42] has published its competitive results, produced by a superior pre-trained model that has been fine-tuned with large-scale data across meticulously defined task categories. Additionally, Emu Edit introduces an instruction-based image-editing benchmark. Each entry in this benchmark includes an input image, a set of instructions, captions for both the input and output images, and an edited result generated by the Emu Edit model. This benchmark encompasses seven distinct tasks: style editing, object removal, object addition, background alteration, color/texture modification, local adjustments, and comprehensive transformations. In our study, we concentrate exclusively on the style editing task, which comprises a total of 226 samples after excluding one sample with blank input and output images.
4.2.2 Evaluation Metrics.
We use CLIP-based metrics as follows: 1) directional score [12] (CLIPdir) measures the alignment of 2 directions, input-output caption change direction and input-output image change direction in CLIP space. The input-output captions are provided in Emu-Edit Benchmark. 2) image similarity (CLIPimg) measures the similarity between the input and edited images indicating the number of changes. 3) output similarity (CLIPout) measures the similarity between the edited image and the output caption.
We also engage crowd workers to assess the editing outcomes. Three human rating scores are presented: the success rate of following the instructions (SRins), the content preservation rate (SRcont), and the user preference rate (UPR), which compares our method against other baseline approaches.
4.2.3 Quantitative Comparisons.
@@ 4.2 Text-Based Style Editing
As displayed in Tab. 1, our method achieves competitive CLIPdir and CLIPout scores in comparison to the previously mentioned baselines. Emu-Edit achieves significantly higher CLIPimg and SRcont scores than others, which means the outputs and the original images are more similar in both CLIP space and human views. It’s essential to measure content-preservation, but an extremely high score might come from insufficient conveying of the style features. We will explain more in qualitative comparison.
Overall, our approach attains the highest instruction-following rate SRins at 55.44% and an elevated user preference rate UPR of 58.85%, outperforming the baseline methods. These figures indicate that our model not only has the highest rate of accurate edits but is also most favored by users.

### v2

@@ 1 Introduction
To tackle these challenges, we introduce StyleBooth, a unified method for text and exemplar-based style editing that accommodates hybrid editing instructions combining written text and exemplar images. Concretely, we devise a unified conditioning schema for the diffusion model, using the tokens "<style>" and "<image>" as identifiers to indicate arbitrary style and exemplar style images. The text instruction, which substitutes the "<style>" token with specified style words, is processed through the text encoder. Additionally, the exemplar image is encoded by the image encoder, and the image feature is aligned with the encoded text feature in a unified hidden space, forming the multimodal input for the diffusion model. Merging the flexibility of natural language with the explicitness of exemplar images enables the exploration of more creative and personalized styles, such as style interpolation and style composition. We present StyleBooth models learning from content-strong-aligned and high-resolution image pairs, which is capable of generating high content fidelity and quality edited images. Samples of edited results are shown in  Fig. 1 and we also present a high-quality style editing dataset, constructed through an effective pipeline combined iterative edit tuning and usability filtering. This strategy is easy to generalize to any other image editing tasks.
@@ 2 Related work
Image style transfer. Image style transfer has been widely studied by researchers. It involves the process of applying stylistic representations to content images in order to generate images with specific styles. Style transfer based on deep neural networks  [13, 14] typically employs separating and recombining the content and style of an image to achieve a simple transfer of style, but there is a problem with the precision of the style representation. There are also works based on GANs  [21, 52, 9, 25] that, through training generators and discriminators, apply the representations of style images while preserving the features of content images as much as possible. However, these methods lack sufficient adaptability and the trained models need to correspond to the style, which limits the possibility of more extensive transformation. Recently, diffusion-based image generation has developed rapidly and has also been applied to style transfer tasks. These methods [50, 8, 42, 46] often require users to provide only a few or even just one style image to achieve a relatively universal style transfer. In this process, generating high-quality transferred images while maintaining the semantic information of the content image remains a challenge. Our StyleBooth builds upon this foundation and can continuously generate high-quality transferred images while preserving the semantic information of the content image.
@@
Text-based image editing. Image editing is a common application in practical scenarios, where it can achieve customized image generation through external guidance. Many pre-trained generative models based on diffusion models [19, 44, 45, 33, 37] can generate corresponding images according to text, but it is difficult to finely edit images according to specific needs while keeping the main content unchanged. Recent works edit global semantic information of images based on text guidance [3, 16, 31]. However, they still require detailed descriptions or corrections of the target images. Text-based image editing offers great convenience, as it only requires describing the specific changes to be made to the image, such as "Let it be in the style of Picasso". InstructPix2Pix [4] is trained with instructional data generated by GPT-3 [5] and Prompt-to-Prompt [16]. MagicBrush [47] builds an image editing dataset with manually annotated instructions using online image editing tools. Emu Edit [41] proposes a multi-task image editing model and trains it on image editing and recognition data. Although these works have achieved improvements, high-quality data and precise execution of editing instructions remain significant challenges.
@@
Exemplar-based style editing. Exemplar-based image editing typically uses one or several image exemplars as references for generations. InST [50] introduces an attention-based textual inversion method that inverts a painting into corresponding text embeddings to guide the model in creating images with a specific artistic appearance. StyleDrop [42] employs an adapter fine-tuning approach to efficiently train models corresponding to the reference images. RAVIL [51] modifies the denoising inference chain of diffusion models to align with the real image inversion chain, generating diversified and high-quality variants of real-world images. The aforementioned methods only support editing with exemplar images, but for more detailed guidance, our StyleBooth can generate highly creative images based on any image exemplar and text instruction while maintaining the consistency of the identity of the images to be edited.
@@
3 StyleBooth
In this section, we present a unified text and exemplar-based style editing method namely StyleBooth. For styles easy to be explained in text, the text-based method is capable of following text instructions. As for those hard to write accurately, one may use an exemplar-based method providing the target style by uploading a reference image. StyleBooth has both advantages. StyleBooth unifies text instruction and exemplar image into a multimodal instruction. Our model is a latent diffusion model conditioned by this multimodal instruction together with an image to be edited as illustrated in Fig. 2. High-quality data is also essential to style editing tasks. We introduce a paired image constructing strategy based on iterative style-destyle editing and usability control. Following this methodology, we present a high-quality dataset supporting text and exemplar-based style editing.
3.1 Multimodal Instruction Editing
Preliminaries. Diffusion models[19] are trained to sample image $x$ from Gaussian noise through a reversed diffusion process, which is a sequence of denoising operations of predicting the added noise $\epsilon\sim\mathcal{N}(0,1)$ at a specific timestamp $t$. To improve the efficiency, latent diffusion[37] utilizes a pre-trained variational autoencoder[26] $\mathcal{E}$ to encode $x$ into latent space $z=\mathcal{E}(x)$ and operates denoising in the compressed latent space. InstructPix2Pix[4], an image editing model, tuned from a pre-trained T2I latent diffusion model $\epsilon_{\theta}$ conditioned by instruction text $c_{T}$ and original image $c_{I}$. $c_{T}$ is encoded by pre-trained CLIP[35] text encoder into text features $h_{T}=\mathcal{C}_{T}(c_{T})$, while $c_{I}$ is encoded by a pre-trained variational autoencoder into latent. The training objective is formulated as follows:
@@ 3.1 Multimodal Instruction Editing
Multimodal instruction. To support text-based and exemplar-based style editing at the same time, our model supports text-only, image-only, and text-image instructions, as illustrated in Fig. 2. Input images are treated as an undetermined identifier "<image>" in the text encoder. We encode the style exemplar image $c_{E}$ using CLIP image encoder $\mathcal{C}_{I}$. The patch features after the last Transformer layer are considered as exemplar image features. We introduce a trainable convolutional alignment layer $W$ to map image features into text feature space:
@@
Textual instruction templates for multimodal inputs. For diverse, comprehensible, and grammar-correct style editing instructions, we leverage the Turbo 128k version of GPT4[34] to generate editing instructions. Instead of complete sentences, we ask LLM for literal templates composed of normal text and identifier tokens. These identifiers represent the undetermined target styles. We write examples to make sure LLM to precisely interpret our intent. For text-based style editing, we give an instruction example "Let this image be in the style of <style>". For exemplar-based style editing, we use "Let this image be in the style of <image>" as an instruction example. Notice that we use "<style>" and "<image>" as the identifiers of arbitrary styles in types of text and exemplar. Adequate templates are generated for each task for training. During inference, input instructions can be varied and unnecessarily limited to the training instructions.
Text-based and exemplar-based style editing. Text-based methods perform editing following text-only instructions. During training, instruction templates are randomly selected and the identifier "<style>" is replaced by the actual style name. We follow the methodology of InstructPix2Pix and fine-tune from their pre-trained model. Supervised by the paired images, the model learns to add the specific style following text-only instructions.
Style of exemplar image and that of target image must be well-aligned during training phase. We search for the most appropriate exemplars from other target images in the same style subset. Within a style subset, we first calculate CLIP score for every target image with other in-class stylized images. Images with the highest score are the most similar ones and they are the most appropriate candidates as exemplar-style images. During training, the final exemplar will be randomly selected from the top 10 most similar images. An instruction template is randomly selected and encoded into text features by $\mathcal{C_{T}}$, where the identifier "<image>" is treated as an unknown identifier at the beginning. Composed of exemplar image features, the original textual instruction is transformed into a multimodal instruction $h$. The training objective is changed into:
@@
Scale weighting mechanism for compositional style editing. As shown in Fig. 2, in order to balance the presence of visual elements, we introduce a Scale Weighting Mechanism. For $n$ textual or visual style representations, we assign a series of multiplicative scale factors $\alpha_{1},\alpha_{2},\ldots,\alpha_{n}$. We find the Scale Weighting Mechanism effective in controlling the style presence ratio in the final results.
@@
Figure 4: Generated samples of the intermediate and final image pairs during Iterative Style-Destyle Tuning and Editing. During iterations, image quality gets higher while key style features are gradually wiped off in the de-styled images. We show the style images and de-style results generated in 1st and 2nd de-styled phase and a plain image and results generated in 1st stylize phase.
@@ 3.2 High-Quality Style Editing Dataset
Generating style and plain images. Intuitively, images in specific styles can be generated by any T2I model as long as the style-related keywords to the text prompt. We follow the same route but design text prompts carefully through prompt expansion provided by Fooocus[11] to avoid the style of generated image misaligned with our prompts. For each of the styles, prompts are modified by the style format before being sent to T2I model. We accept 67 prompt formats and 217 diverse content prompts, ending up with 67 different styles and 217 images per style. This batch of stylized images is marked as $BatchA$. Next we generate $BatchB$ consisting of plain images. For diversity and quality of generated images, we randomly select image captions from the LAION Art[40] dataset as prompts. Unlike style image generation, no specific style prompt expansion is used.
@@
Vanilla de-style and usability filtering. We train a vanilla version of image editing model following the methodology of Instruct-Pix2Pix[4]. This model has a basic I2I editing ability but the generated data quality falls far short of our needs. Starting from $BatchA$, we use the image editing model to de-style them into plain photographic images generating the corresponding $BatchA^{1}$. As shown in the second column of Fig. 4, after the first de-style operation, the content structure is preserved yet most of these de-styled images contain features of the original style. We then employ a CLIP-based[35] metric to filter out image pairs that are too similar, indicating no significant style change, and that are overly different, which could imply potential content distortion. By establishing upper and lower thresholds for CLIP scores, this filtering process effectively eliminates the majority of evident failures produced by the image editing model.
@@
Iterative style-destyle tuning and editing. To further strengthen the pre-trained I2I model, we take filtered image pairs as training data and fine-tune it using Style Tuner to learn style-adding ability. Now, the $BatchA^{1}$ are used as inputs while the style images in $BatchA$ as targets. Efficient tuning methods[20, 23, 15, 24] show strong few-shot learning capability. Here, we add efficient tuners called Style Tuners to the basic I2I model and tune different styles separately which means each tuner is only responsible for one specific style. After tuning, the base model with tuners will edit $BatchB$ images to add the styles they have learned. Then we get $BatchB^{1}$. Some of the results are shown in the last column of Fig. 4. The style-adding results is frustrating and styles in $BatchB^{1}$ is still not that accurate. Again, $BatchB$ and $BatchB^{1}$ images are measured and filtered by CLIP-based metric to be the next round training data. Note that we always use the high-quality images in $BatchA$ or $BatchB$ as the training targets. In this round, we tune I2I model to be a style removal model using De-style Tuner, which means images in $BatchB^{1}$ are used as inputs and image in $BatchB$ as targets. Still, one tuner is trained to remove one type of style. The tuned I2I model edits the stylized image in $BatchA$ and geneates $BatchA^{2}$ which are shown in the third row of Fig. 4. The image quality in $BatchA^{2}$ is improved dramatically. We also use CLIP scores to filter out unqualified image pairs. After $n$ iterations, $BatchA^{n}$ are generated and image pairs of $BatchA$ and $BatchA^{n}$ are filtered. The style-diverse, high-quality and content-preserved stylized/plain image pairs are finally observed. We compare the number of images those passing the CLIP filter. Comparing to the results of vanilla de-style, the average usability rate is increased dramatically from 38.11% to 79.91% by 41.80%, which shows that our Iterative Style-Destyle Editing is effective to improve paired images’ quality.
@@ 4.1 Experimental setup
We conduct our data generation and experiments on an open-source project SCEPTER[39], integrating many styles from Fooocus[11] and providing an easy implementation of efficient tuning methods. We use SDXL[1] to generate $BatchA$ and $BatchB$. LoRA[20] with rank 256 are accepted as Style Tuners and De-style Tuners. The editing-tuning iterations stop at the end of the 2nd de-style process. So the final training pairs are $BatchA^{2}$ and $BatchA$. The trainable $W$ is configured as a $6\times 6$ convolutional layer with stride 4, mapping a $14\times 14$ visual features into 9.
@@
4.2 Text-Based Style Editing
Baselines and benchmarks. We compare our model with 3 state-of-the-art text-based image editing methods. Both InstructPix2Pix[4], Magic Brush[47] and The Emu Edit[41] present machine-generated image editing instruction-tuning datasets involving triplets of the original image, edited image, and text instruction. Magic Brush mainly focuses on local editing tasks while InstructPix2Pix and Emu Edit are more generalized. Additionally, Emu Edit introduces an instruction-based image-editing benchmark. Each entry in this benchmark includes an input image, a set of instructions, captions for both the input and output images, and an edited result generated by the Emu Edit model. This benchmark encompasses seven distinct tasks: style editing, object removal, object addition, background alteration, color/texture modification, local adjustments, and comprehensive transformations. In our study, we concentrate exclusively on the style editing task, which comprises a total of 226 samples after excluding one sample with blank input and output images.
Evaluation metrics. We use the following CLIP-based metrics: (1) Directional score(CLIPdir)[12] measures the alignment of 2 directions, input-output caption change direction and input-output image change direction in CLIP space. The input-output captions are provided in Emu-Edit Benchmark. (2) Image similarity(CLIPimg) measures the similarity between the input and edited images indicating the number of changes. (3) Output similarity(CLIPout) measures the similarity between the edited image and the output caption. To assess editing outputs under human standards, three human rating scores are presented: (1) Instruction success rates(SRins) estimates whether output styles are correct and match instructions. (2) Content success rate(SRcont) estimates the content consistency. (3) User preference rate(UPR) compares our method against other baseline approaches.
Quantitative comparisons. As displayed in Tab. 1, our method achieves competitive CLIPdir and CLIPout scores in comparison to the previously mentioned baselines. Emu-Edit achieves significantly higher CLIPimg and SRcont scores than others, which means the outputs and the original images are more similar in both CLIP space and human views. It’s essential to measure content-preservation, but an extremely high score might come from insufficient conveying of the style features. We will explain more in qualitative comparison. Overall, our approach attains the highest instruction-following rate SRins at 55.44% and an elevated user preference rate UPR of 58.85%, outperforming the baseline methods. These figures indicate that our model not only has the highest rate of accurate edits but is also most favored by users.